OpenStudy (anonymous):

Will fan and medal no straight answers Complete the standard form of the equation –4y = 3 – 2x. _x+_y=3 (Basically you just fill in the blanks but dont give me straight answers please explain it)

OpenStudy (anonymous):

@baby_bear69 can you halp pls with your math smartness

OpenStudy (goose.):

Well, -4y would equal 3 correct?

OpenStudy (anonymous):

I just don't understand what they mean by complete the standard form. And yeah

OpenStudy (baby_bear69):

\[-4y=3-2x\]

OpenStudy (baby_bear69):

add 2 x to both sides

OpenStudy (baby_bear69):

-4y +2x +3

OpenStudy (baby_bear69):

=3***

OpenStudy (baby_bear69):

i guess thats what they want

OpenStudy (anonymous):

I'm still a bit confused im so sorry lol

OpenStudy (baby_bear69):

its called isolating the variable to get a constant (regular number) on one side and the variables on the other

OpenStudy (anonymous):

Okay so what's the first step sorry again for being stupid xP

OpenStudy (jhannybean):

standard form of an equation: \(\sf Ax~ +~By=C\)

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):

seems like you got it.

OpenStudy (anonymous):

I hope so

OpenStudy (anonymous):

never mind I think I got it now

OpenStudy (jhannybean):

Ok.

OpenStudy (anonymous):

Would -4x+(-2x)=3 be correct

OpenStudy (jhannybean):

-4y, right?

OpenStudy (anonymous):

yeah yeah that's what I meant sorry

OpenStudy (jhannybean):

Yes, and we can simplify that to say -4y - 2x = 3, because a + times a - gives you a -

OpenStudy (anonymous):

That was so simple I feel so stupid now, lol thanks